OpenStudy (anonymous):

"Which Term of 127, 122, 117, 112,... is -68 ?" Its an Arithmetic Problem. Alg 2. I know the A sub n formula and the S sub n formula i just don't understand how to apply?

OpenStudy (mkapad01):

Un= 127-5(n-1) plug in -68 for n

OpenStudy (mimi_x3):

Ok so the \(A_n\) formula is just the formula that states the value of a specific term in the sequence so \[A_n=A_1+(n-1)d\] So where A_1 is our first term in the sequence and d stands for the difference btwn each term in the sequence. n stands for the sequence number

OpenStudy (mimi_x3):

So A_1 in our case is 127 .. do you agree??? Thats our first term in the sequence

OpenStudy (anonymous):

okay understood but when i plug in 127 for N it gives me An = - 208

OpenStudy (mimi_x3):

Why u plugging in 127 for n????

OpenStudy (anonymous):

i mean -68* my bad

OpenStudy (mimi_x3):

So no n doesnt represent -68 ... -68 =A_n We wanna find out what our n is So basically we know the value of the term but what term number is it with regard to the sequence???? SO is it our 30th term in the sequence?? Or is it our 40th term in our sequence??? We dont know so we have to use the formula A_n to find out what our n equals to

OpenStudy (mkapad01):

just to mention the (a sub n) could be anything, it could be u of n or t of n. Just a variable

OpenStudy (mimi_x3):

yes thats true ... Im doing a bad job of explaining here -.-

OpenStudy (mkapad01):

a sub n is like y

OpenStudy (mkapad01):

find the equation, and plug -67 into An

OpenStudy (mkapad01):

The equation I stated in the beginning should be right

OpenStudy (anonymous):

okay so basically i just have to plug in each number ( 122, 127, 117, 112) as N until one equals -68 ?

OpenStudy (mimi_x3):

Hey lets start from the beggining of sequences ok

OpenStudy (mkapad01):

no, plug into An, and solve for n

OpenStudy (mkapad01):

@Mimi_x3 got this :)

OpenStudy (mimi_x3):

So we have an arithmetic sequence that starts with 127, 122, 117, 112 .... So our first term is \(A_1=127\) Our second term of the sequence is \(A_2=122\) Our third term of the sequence is \( A_3=117\) Our fourth term is \(A_4=112\) And our sequence just continues on forever How does this arithmetic sequence work???? Well basically a certain number is subtracted/added from every term and this is known as the difference How do we find the difference? We subtract the second term from the first term \( \text{ Difference } =A_2-A_1=122-127=-5\) So every -5 is subtracted from the term to get the next term in the sequence 122-127=-5 117-122=-5 112-117=-5 As you can see The way figure out this sequence is by subtracting 5 K so we have this formula: \[A_n=A_1+(n-1)d\] \(A_n \)is the value of the term n represents the sequence number or like what term in the sequence it is and d represents the difference So the question is saying -68 is one of the terms in the sequence but what number term is it???? So using the formula where \( A_n=-68\) and d=-5 find plug in these values and solve for n

OpenStudy (dan815):

tsk tsk

OpenStudy (mimi_x3):

Dan lmao

OpenStudy (mimi_x3):

just fukin medal me

OpenStudy (mimi_x3):

@Kainui Pls medal me as Dan refuses to

OpenStudy (dan815):

okay listen to me buddy i want to show u something cool

OpenStudy (mimi_x3):

lmao this buddy doesnt even get the whole sequence thing ..... Ik what u wanna do SHow how to derive the A_n formula lol

OpenStudy (anonymous):

well haha thanks!

OpenStudy (dan815):